Thirty-two small communities in Connecticut (population near 10,000 each) gave an average of x = 138.5 reported cases of larceny per year. Assume that σ is known to be 40.5 cases per year.

(a) Find a 90% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)

lower limit    
upper limit    
margin of error    


(b) Find a 95% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)

lower limit    
upper limit    
margin of error    


(c) Find a 99% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)

lower limit    
upper limit    
margin of error    
Answer #1

Part a)

Confidence Interval :-
\bar{X} \pm Z_{\alpha /2} \sigma /\sqrt{n}
Z_{\alpha /2} = Z_{ 0.1 /2} = 1.64
138.5 ± Zo.1/2 * 40.5/V32
Lower Limit = 138.5 - Z0.1/2 40.5/V32
Lower Limit = 126.7237
Upper Limit = 138.5 + Zo.i/2 40.5/V32
Upper Limit = 150.2763
90% Confidence interval is ( 126.7 , 150.3 )

Margin of Error =   Z0.1/2 40.5/V32- 11.8

Part b)

Confidence Interval :-
\bar{X} \pm Z_{\alpha /2} \sigma /\sqrt{n}
a/2-20.05/2 -1.96
20.05/2 * 40.5/ V32 138.5
Lower Limit = 138.5 - Zo.05/2 40.5/v32
Lower Limit = 124.4677
Upper Limit = 138.5 + Zo.05/2 40.5/V32
Upper Limit = 152.5323
95% Confidence interval is ( 124.5 , 152.5 )


Margin of Error =   Zo.1/2 40.5/V32- 14.0

Part c)

Confidence Interval :-
\bar{X} \pm Z_{\alpha /2} \sigma /\sqrt{n}
Z_{\alpha /2} = Z_{ 0.01 /2} = 2.58
20.01/2 * 40.5/ V32 138.5
Lower Limit = 138.5 - Z_{ 0.01/2} \;40.5/\sqrt{ 32}
Lower Limit = 120.0585
Upper Limit = 138.5 Zo.o1/2 40.5/V32
Upper Limit = 156.9415
99% Confidence interval is ( 120.1 , 156.9 )

Margin of Error =   Z_{ 0.1/2} \;40.5/\sqrt{ 32} = 18.4